OpenStudy (anonymous):

a4(a-1)(a3) A.a7 B.a6 C.a-12

OpenStudy (campbell_st):

is your question \[a^4\times a^{-1}\times a^3\] here is a help.... the index law for multiplying the same base \[x^a \times x^b = x^{a + b}\] use the law to get your answer.

OpenStudy (anonymous):

thank you very much